# The integral Chow ring of $\mathcal{R}_{2}$

### v1

@@ 3.4 Proof of Theorem 4
Which, after using (23), the substitution $h=\beta_{1}+\gamma$, and the relations coming from $\mathrm{CH}^{*}(BG)$, yield $-2\beta_{1}$, $8\beta_{2}-3\beta_{1}(\beta_{1}+\gamma)$ and $-8\beta_{2}$, respectively, and linear combinations of these give the desired elements. ∎
@@
In particular, $\alpha\cdot\mathrm{mult}(s_{a}^{a^{\prime}},s_{4-a}^{b^{\prime}})=0$ whenever $(\beta_{1}+\gamma)\mid\alpha$, or $a^{\prime}+b^{\prime}\geq 1$ and $2\mid\alpha$, or $a^{\prime}+b^{\prime}\geq 3$.
@@
We have to pushforward all the $s_{r}^{i}$ classes along maps $p_{i,j}$. Note that $p_{2,2}$ and $p_{3,2}$ factor through $p_{1,1}$, and so there is no need to consider them.
@@
Pushforward along $p_{4}$: This is just the fundamental classes of $\{X^{4},Y^{4}\}$, $\{X^{3}Y,XY^{3}\}$ and $\{X^{2}Y^{2}\}$, which we calculated in Lemma 34, and are easily seen to be 00. ∎
@@ 4 Interpretation of the generators
In the proof of the above proposition we will refer to [Vis98, Proposition 3.1]. However, differently from what is claimed there, with the notation of that paper, the action of $\mathrm{GL}_{2}$ induced on the space $X$ by the isomorphism with $Y$ and the action on $Y$ identifies $X$ with the open locus in $\mathrm{Sym}^{6}(V)\otimes\mathrm{det}(V)^{\otimes 2}$ where the polynomial has no repeated roots. In the paper, it was instead claimed that $X\subseteq\mathrm{Sym}^{6}(V^{\vee})\otimes\mathrm{det}(V)^{\otimes 2}$.
Proof.
It is enough to observe that we have a commutative square
${\mathcal{X}}$${\mathcal{R}_{2}}$${\bigg{[}\frac{X}{\mathrm{GL}_{2}}\bigg{]}}$${\mathcal{M}_{2}}$$\scriptstyle{\alpha}$$\scriptstyle{\varphi}$
where the bottom row is the map in [Vis98] and the map $\varphi$ is defined as follows.
Write $V=\mathbb{C}^{2}$ as vector spaces and let $e_{1}=(1,0)$ and $e_{2}=(0,1)$ be the canonical basis and $X,Y\in V^{\vee}$ be the dual basis. Then given
$G=\sum_{i_{1}+i_{2}=6}a_{i_{1},i_{2}}X^{i_{1}}Y^{i_{2}}\in\mathrm{Sym}^{6}(V^{\vee})$
we obtain an element $\widetilde{G}\in\mathrm{Sym}^{6}(V)$ as follows:
$\widetilde{G}=\sum_{i_{1}+i_{2}=6}a_{i_{1},i_{2}}e_{1}^{i_{1}}e_{2}^{i_{2}}.$
Then
$\varphi(F):=\widetilde{XYF}$
and $G\hookrightarrow\mathrm{GL}_{2}$ is given by
$A\mapsto(A^{-1})^{t}$ (31)
By [Vis98, Proposition 3.1] we know that the pullback of the Hodge bundle to $\bigg{[}\frac{X}{\mathrm{GL}_{2}}\bigg{]}$ is $V_{G}$ and thus, by (31), the further pullback to $\mathcal{X}$ is $V_{G}^{\vee}$ as desired. ∎
@@
${D=\sigma_{0}(U)\sqcup\sigma_{\infty}(U)}$${C=\underline{\mathrm{Spec}}_{\mathbb{P}^{1}_{U}}(\mathcal{O}_{\mathbb{P}^{1}_{U}}\oplus\mathcal{O}_{\mathbb{P}^{1}_{U}}(-3))}$${\mathbb{P}^{1}_{U}=U\times\mathbb{P}(V)}$${U}$$\scriptstyle{\delta}$$\scriptstyle{q}$$\scriptstyle{\pi}$$\scriptstyle{\rho}$
@@
$\bigg{[}\frac{D}{G}\bigg{]}\to\mathcal{R}_{2}$
and have the following lemma.
@@
The pushforward of the structure sheaf $\mathcal{O}_{D}$ under $\delta$ is clearly $\mathcal{O}_{U}\otimes A$ where $A$ is the two dimensional representation of $G$ arising from the representation of $\mathbb{Z}/(2)$ where $-1$ switches the two entries of the vector. Clearly, $\mathrm{det}(A)=\Gamma$ and the lemma follows. ∎

### v2

@@ 3.4 Proof of Theorem 4
Which, after using (23), the substitution $h=\beta_{1}+\gamma$, and the relations coming from $\mathrm{CH}^{*}(BG)$, yield $2\beta_{1}$, $8\beta_{2}-3\beta_{1}(\beta_{1}+\gamma)$ and $8\beta_{2}$, respectively, and linear combinations of these give the desired elements. ∎
@@
An immediate but useful consequence is the following.
Remark 37.
We have $\alpha\cdot\mathrm{mult}(s_{a}^{a^{\prime}},s_{4-a}^{b^{\prime}})=0$ whenever $(\beta_{1}+\gamma)\mid\alpha$, or $a^{\prime}+b^{\prime}\geq 1$ and $2\mid\alpha$, or $a^{\prime}+b^{\prime}\geq 3$.
@@
We have to pushforward all the $s_{r}^{i}$ classes along maps $p_{i,j}$. Note that $p_{2,2}$ and $p_{3,2}$ factor through $p_{1,1}$, and so there is no need to consider them. Remark 37 will be used repeatedly and without further mention.
@@
Pushforward along $p_{4}$: This is just the fundamental classes of $\{X^{4},Y^{4}\}$, $\{X^{3}Y,XY^{3}\}$ and $\{X^{2}Y^{2}\}$, which we calculated in Lemma 35, and are easily seen to be 0. ∎
@@ 4 Interpretation of the generators
In the proof of the above proposition we will refer to [Vis98, Proposition 3.1]. However, with the notation of that paper, there is an error in the description of the $\mathrm{GL}_{2}$-action on the space $X$. It is incorrectly stated that this action, induced by the isomorphism of $X$ with the $\mathrm{GL}_{2}$-equivariant variety $Y$, identifies $X$ as an open subset of $\mathrm{Sym}^{6}(V^{\vee})\otimes\det(V)^{\otimes 2}$. Instead, $X$ is identified with the open subset of $\mathrm{Sym}^{6}(V)\otimes\det(V^{\vee})^{\otimes 2}$ consisting of polynomials with no repeated roots.
Base changing over the morphism
$B\mathrm{GL}_{2}\to B\mathrm{GL}_{2}$
given by $A\mapsto(A^{-1})^{t}$, we get an isomorphism
$\mathcal{M}_{2}\cong\left[X^{\prime}/\operatorname{GL}_{2}\right],$ (31)
where $X^{\prime}\subset\mathrm{Sym}^{6}(V^{\vee})\otimes\det(V)^{\otimes 2}$ is the locus with no repeated roots, and under this new isomorphism, the Hodge bundle is the pullback of $V_{\operatorname{GL}_{2}}^{\vee}$ from $B\operatorname{GL}_{2}$.
Proof of Proposition 40.
The morphism
$\chi:\operatorname{Sym}^{4}(V^{\vee})\otimes\det(V)\otimes\Gamma$ $\longrightarrow$ $\,\mathrm{Sym}^{6}(V^{\vee})\otimes\det(V)^{\otimes 2}$
$F$ $\mapsto$ $XYF$
is equivariant with respect to the inclusion $\iota:G\to\operatorname{GL}_{2}$, and so we have a commutative diagram
$\mathcal{X}$$\left[X^{\prime}/\operatorname{GL}_{2}\right]$$\mathcal{R}_{2}$$\mathcal{M}_{2},$$\scriptstyle{\alpha}$$\scriptstyle{\tilde{\chi}}$(31)
where the bottom arrow is the forgetful morphism. Therefore, the proposition follows because the pullback of the Hodge bundle under (31) is $V_{\operatorname{GL}_{2}}^{\vee}$, and $\iota^{*}V_{\operatorname{GL_{2}}}=V_{G}$. ∎
@@
[…]
@@
$\bigg[\frac{D}{G}\bigg]\to\mathcal{R}_{2},$
and we have the following lemma.
@@
The pushforward of the structure sheaf $\mathcal{O}_{D}$ under $\delta$ is clearly $\mathcal{O}_{U}\otimes A$ where $A$ is the two dimensional representation of $G$ arising from the representation of $\mathbb{Z}/(2)$ where 1-1 switches the two entries of the vector. Clearly, $\mathrm{det}(A)=\Gamma$ and the lemma follows. ∎